OpenStudy (anonymous):

The formula for determining interest compounded monthly is A = P(1 + r/12 )^12t, where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $1,000 at an interest rate of 10% for 3 years, while Jenny invests $1,000 at an interest rate of 5% for 6 years. Determine the amount of return gained by Jimmy and Jenny. In complete sentences, summarize your results.

OpenStudy (anonymous):

since it was compounded annually just sub in the values: A= (1000)(1.1)^3, AND A= (1000)(1.05)^6, compare the two values

OpenStudy (anonymous):

what does that mean ?

OpenStudy (anonymous):

what part are you confused about

OpenStudy (anonymous):

how did you get your numbers like 1.1

OpenStudy (anonymous):

1+r = 1 +10% =1+0.1 =1.1

OpenStudy (anonymous):

A= (1000) (1 + 10/12) ^12(3) ?

OpenStudy (anonymous):

is it like that ?

OpenStudy (anonymous):

no, as you don't need to divide by 12, as everything is compounded annually

OpenStudy (anonymous):

okay ! so what do I do now with A= (1000)(1.1)^3, AND A= (1000)(1.05)^6

OpenStudy (anonymous):

the formula for compound interest is a=p(1+r)^n, where a= interest earned, p= principle, basically what you put in, r= the rate of the interest in decimal place and n= no of years term is invested

OpenStudy (anonymous):

yeah that is true, sorry misread the question

OpenStudy (anonymous):

okay ! so what do I do now with A= (1000)(1.1)^3, AND A= (1000)(1.05)^6

OpenStudy (kropot72):

The question makes it clear the the interest is compounded monthly. The annual interest rate must be expressed as a decimal to use it in the formula. So 10% becomes 0.1 and the monthly interest rate is 0.1/12. Jimmy's return is found by the calculation \[A=1000(1+\frac{0.1}{12})^{(3\times12)}\]

OpenStudy (kropot72):

Note that the formula given in the question must be used.

OpenStudy (anonymous):

ooh okay ! & now what do I do ?

OpenStudy (kropot72):

First calculate the sum of the terms inside the brackets, then raise that sum to the power of 36. Next multiply the result by 1000.

OpenStudy (anonymous):

1,348.18 ?

OpenStudy (kropot72):

Good work! .Your result is correct

OpenStudy (anonymous):

okay so when I add both numbers, I get $18,900 .41 + $1,348.18 = $20,248.58 !

OpenStudy (kropot72):

Where did you get $18,900.41 from???

OpenStudy (anonymous):

the second one ! A= (1000)(1 + 0.5/12)^72 = 18,900 .41

OpenStudy (anonymous):

right ?

OpenStudy (kropot72):

Not really. Jenny's return is found from \[A=1000(1+\frac{0.05}{12})^{72}\] The annual interest rate of 5% is 0.05 expressed as a decimal/

OpenStudy (anonymous):

so this ? $ 1,348.18 + $1,348.18 = $2,697.19

OpenStudy (kropot72):

My calculation for Jenny's return came to $1349.02. The summary of the results needs to make the following point: Jenney's investment at half the interest rate and double the investment term of Jimmy's made only a slightly higher return. Obviously Jimmy made the better investment.

OpenStudy (anonymous):

can you type how you did that ?

OpenStudy (kropot72):

Item Jenny Jimmy Investment $1000 $1000 Annual rate 5% 10% Investment period 6 years 3 years Return $1349.02 $1348.18